OpenStudy (anonymous):

What is the slope, and what is the y-intercept of the line? y= 5/4x + 9

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

slope = 5/4 y int = 9

OpenStudy (anonymous):

y-int = 9

OpenStudy (alexandervonhumboldt2):

y=ax+b where a is lope and b is y-intercept